OpenStudy (anonymous):

Write the slope-intercept form of the line which passes through (4, 2) with a slope of 1/2.

OpenStudy (anonymous):

y-2=1/2(x-4)

OpenStudy (anonymous):

y=1/2x-2+2=1/2x

OpenStudy (anonymous):

y=1/2x